OpenStudy (anonymous):

give exact and approximate solutions to three decimal place (x-8)^2=16

OpenStudy (anonymous):

(x-8)=sqrt(16) X-8=-4 and +4 x=4 and 12

OpenStudy (mertsj):

Taking the square root of both sides yields x - 8 = 4 or -4. Add eight to both sides and get 12 or 4 so 12.000 and 4.000